OpenStudy (anonymous):

Find the quotient . 6/7 / 9/10 ? help

OpenStudy (anonymous):

6/7 x 10/9

OpenStudy (lgbasallote):

6/7 x 10/9 = 60/63 = 20/21

OpenStudy (anonymous):

u must multiply by the reciprocal